### v2

# Holographic Tensor Networks as Tessellations of Geometry

Abstract
Holographic tensor networks serve as toy models for the Anti-de Sitter/Conformal Field Theory (AdS/CFT) correspondence, capturing many of its essential features in a concrete manner. However, existing holographic tensor network models remain far from a complete theory of quantum gravity. A key obstacle is their discrete structure, which only approximates the semi-classical geometry of gravity in a qualitative sense. In Lin:2024dho , it was shown that a network of partial-entanglement-entropy (PEE) threads, which are bulk geodesics with a specific density distribution, generates a perfect tessellation of AdS space. Moreover, such PEE-network tessellations can be constructed for more general geometries using the Crofton formula. In this paper, we assign a quantum state to each vertex in the PEE network and develop two holographic tensor network models: the factorized PEE tensor network, which takes the form of a tensor product of EPR pairs, and the random PEE tensor network. In both models we reproduce the exact Ryu–Takayanagi formula by showing that the minimal number of cuts along a homologous surface in the network exactly computes the area of this surface.
1 Introduction
The Anti-de Sitter/Conformal Field Theory (AdS/CFT) correspondence Witten:1998qj ; Gubser:1998bc ; Maldacena:1997re , also known as the holographic principle, is a profound framework in theoretical physics that establishes an equivalence between a gravitational theory in a higher-dimensional spacetime and a quantum field theory on its lower-dimensional boundary. It provides a non-perturbative definition of quantum gravity in AdS space in terms of a well-defined quantum field theory without gravity. Moreover, the duality has deep connections to quantum information theory. For example, concepts like entanglement wedge reconstruction (or subregion duality) Czech:2012bh ; Hamilton:2006az ; Morrison:2014jha ; Bousso:2012sj ; Bousso:2012mh ; Hubeny:2012wa ; Wall:2012uf ; Headrick:2014cta ; Jafferis:2015del ; Dong:2016eik ; Cotler:2017erl reveal how quantum information in a bulk region is encoded in the boundary theory, leading to natural error-correcting properties Almheiri_2015 ; Dong:2016eik ; Kamal:2019skn ; Kang:2019dfi ; Akers:2021fut ; Xu:2024xbz ; Crann:2024gkv .
The tensor network approach Swingle:2009bg ; Swingle:2012wq ; Qi:2013caa ; Miyaji:2015fia ; Czech:2015kbp ; Harlow:2016vwg ; Vasseur:2018gfy ; Dong:2018seb ; Hung:2019zsk ; Bao:2019fpq provides a concrete framework to explore the structure of the AdS/CFT correspondence 111see Pastawski:2015qua ; Hayden:2016cfa ; Yang:2015uoa ; Qi:2013caa ; Bao:2019fpq ; Cheng:2022ori ; Dong:2023kyr for an incomplete list of recent progress.. More explicitly, we start with a space filled with tensors representing quantum states on each site in the space, then we contract the “legs” (indices) to connect adjacent tensors thus get a tensor network, which plays the role of the background geometry. The uncontracted legs of tensors in the bulk and the boundary are regarded as the bulk and the boundary degrees of freedom respectively, and the tensor network provides a quantum-error-correcting code that encodes the bulk states into the boundary Hilbert space Dong:2016eik ; Hayden:2016cfa . Many toy models of tensor network to simulate the AdS/CFT correspondence have been proposed (see also Chandra:2023dgq ; Chen:2024unp ; Bao:2024ixc ; Kaya:2025vof ; Akella:2025owv for recent interesting developments), based on different structures of the networks and different quantum states associated to the tensors. One important holographic feature for these toy models is that, the entanglement entropy is given by the minimal number of cuts along a homologous surface in the network, which matches the area of the minimal surface of the celebrated Ryu-Takayanagi (RT) formula Ryu:2006bv ; Hubeny:2007xt ; Casini_2011 ; Lewkowycz:2013nqa ; Faulkner:2013ana ; Engelhardt:2014gca .
Nevertheless, in most of these toy models their structures of the tensor network are usually discrete and only capture the features of the background geometry in a qualitative way. There is a big gap between number of cuts in a discrete network and the area of a surface in a smooth Riemannian manifold. Recently, a continuous network as a perfect tessellation of the background AdS geometry was constructed using the so-called partial-entanglement-entropy (PEE) threads Lin:2023rxc ; Lin:2024dho , which are bulk geodesics with a particular density distribution determined by the PEE structure of the boundary CFT. In such a network, the number of cuts along a surface exactly reproduces the area of this surface. In this work, we will develop two toy models of tensor network based on the PEE threads (or the PEE tensor network for short) in $d$-dimensional Poincaré AdS space. The first one is the factorized PEE tensor network constructed via tensor product of EPR pairs. It enjoys similar properties to the Harlow-Pastawski-Preskill-Yoshida (HaPPY) code Pastawski:2015qua which is built from perfect tensors, and perfectly reproduces the RT formula for spherical regions in the vacuum CFT state. In the second model, following Hayden:2016cfa ; Vasseur:2018gfy ; Jia:2020etj ; Cheng:2022ori ; Dong:2023kyr ; Kaya:2025vof we associate a random state to each bulk site to build the random PEE tensor network, where we can exactly reproduce the RT formula for generic boundary regions.
2 The network of PEE threads
The partial entanglement entropy (PEE) Wen:2018whg ; Wen:2019iyq ; Wen:2020ech ; Han:2019scu ; Han:2021ycp ; Kudler-Flam:2019oru is a measure of entanglement $\mathcal{I}(A,B)$ between any two non-overlapping regions $A$ and $B$ featured by the key property of additivity. It can be determined by the following physical requirements Wen:2019iyq ; Vidal:2014aal :
  • 1.
    Additivity: $\mathcal{I}\left(A,B\cup C\right)=\mathcal{I}\left(A,B\right)+\mathcal{I}\left(A,C\right)$;
  • 2.
    Permutation symmetry: $\mathcal{I}\left(A,B\right)=\mathcal{I}\left(B,A\right)$;
  • 3.
    Normalization: $\mathcal{I}\left(A,\bar{A}\right)=S_{A}$;
  • 4.
    Positivity: $\mathcal{I}\left(A,B\right)>0$;
  • 5.
    Upper bounded: $\mathcal{I}\left(A,B\right)\leq\min\left\{S_{A},S_{B}\right\}$;
  • 6.
    $\mathcal{I}\left(A,B\right)$ should be invariant under local unitary transformations inside $A$ or $B$;
  • 7.
    Symmetry: For any symmetry transformation $\mathcal{T}$ under which $\mathcal{T}A=A^{\prime}$ and $\mathcal{T}B=B^{\prime}$, we have $\mathcal{I}\left(A,B\right)=\mathcal{I}\left(A^{\prime},B^{\prime}\right)$.
Here the regions are all non-overlapping, $A\cup\bar{A}$ makes a pure state and $S_{A}$ denotes the entanglement entropy of $A$. The solutions for these requirements were first studied in Casini_2011 as the mutual information that is additive in some special theories, and then the solutions are understood as the PEE in Wen:2019iyq , which is in general not a mutual information. Later it was gradually realized that, the requirement of normalization should not be imposed for an arbitrary region. In the case of the vacuum CFT state on a plane, the solution only exists if we impose the normalization requirement to spherical regions Wen:2019iyq . See also Vidal:2014aal ; Wen:2018whg ; Wen:2018mev ; Wen:2019iyq ; Kudler-Flam:2019oru for other prescriptions to construct the PEE which satisfy the above requirements, and see Basu:2023wmv ; Rolph:2021nan ; Lin:2022aqf ; Lin:2023orb ; Wen:2021qgx ; Wen:2022jxr ; Wen:2024uwr ; Camargo:2022mme for related discussions on PEE.
Due to the properties of additivity and permutation symmetry, any PEE $\mathcal{I}(A,B)$ can be further decomposed into a class of PEE between two different sites $\mathbf{x}$ and $\mathbf{y}$, or the two-point PEE $\mathcal{I}\left(\mathbf{x},\mathbf{y}\right)$, in the following way Wen:2019iyq ; Lin:2023rxc :
$\mathcal{I}\left(A,B\right)=\int_{A}d\mathbf{x}\int_{B}d\mathbf{y}\>\mathcal{I}\left(\mathbf{x},\mathbf{y}\right).$ (1)
Moreover, in the vacuum CFTd on a plane the two-point PEE can be fully determined by imposing all the above requirements for the spherical regions Lin:2023rxc ; Lin:2024dho , i.e.
$\mathcal{I}\left(\mathbf{x},\mathbf{y}\right)=\frac{c}{6}\frac{2^{\left(d-1\right)}\left(d-1\right)}{\Omega_{d-2}\left|\mathbf{x}-\mathbf{y}\right|^{2\left(d-1\right)}},$ (2)
where $\Omega_{d-2}=2\pi^{\frac{d-1}{2}}/\Gamma\left(\frac{d-1}{2}\right)$ is the area of $\left(d-2\right)$-dimensional unit sphere.
In the context of AdS/CFT, the vacuum CFTd state on a plane is dual to the Poincaré AdSd+1 space. The two-point PEE $\mathcal{I}\left(\mathbf{x},\mathbf{y}\right)$ was also proposed to have a natural holographic dual Lin:2023rxc ; Lin:2024dho , which is the density of the bulk geodesics connecting the boundary sites $\mathbf{x}$ and $\mathbf{y}$. The bulk geodesics whose density are determined in such a way are denoted as the PEE threads. Let us set that each PEE thread gives unit contribution to the PEE, then (1) captures the number of PEE threads connecting the two boundary regions $A$ and $B$. The collection of all the PEE threads forms a continuous, dense network permeating the entire time slice of the Poincaré AdSd+1, which we call the PEE network (see an example of $d=2$ in Fig.1). Remarkably, it was found that Lin:2023rxc ; Lin:2024dho the PEE network is a tessellation of the static AdS space in the following sense:
  • Given an arbitrary (d1)(d-1)-dimensional surface $\Sigma$ in a time slice of AdSd+1, the density of intersections between $\Sigma$ and the bulk PEE network everywhere on $\Sigma$ is given by a constant $\frac{1}{4G}$, where $G=\frac{3}{2c}$ is the Newton’s constant.
Then for any bulk surface $\Sigma$, the area can be reproduced by counting the number $\mathcal{N}\left(\Sigma\right)$ of intersections between $\Sigma$ and the PEE network,
$\frac{\text{Area}(\Sigma)}{4G}=\mathcal{N}\left(\Sigma\right)=\frac{1}{2}\int_{\partial\mathcal{M}}d\mathbf{x}\int_{\partial\mathcal{M}}d\mathbf{y}\>\omega_{\Sigma}\left(\mathbf{x},\mathbf{y}\right)\mathcal{I}\left(\mathbf{x},\mathbf{y}\right).$ (3)
Here $\mathcal{M}$ is a constant time slice in the Poincaré AdSd+1, while $\partial\mathcal{M}$ is the boundary where the vacuum CFT state lives, and $\omega_{\Sigma}\left(\mathbf{x},\mathbf{y}\right)$ represents of the number of intersections between $\Sigma$ and the particular PEE thread connecting boundary sites $\mathbf{x}$ and $\mathbf{y}$.
[…]
$\text{Area}\left(\Sigma\right)=\frac{1}{2}\frac{d-1}{\Omega_{d-2}}\int_{\text{Geod}\left(M\right)}\#\left(\Gamma\cap\Sigma\right)d\Gamma.$ (4)
Here, the integrand $\left(\Gamma\cap\Sigma\right)$ represents the number of intersections between any geodesic $\Gamma$ in $\text{Geod}\left(M\right)$ and $\Sigma$, and the Kinematic measure $d\Gamma$ specifies the density of geodesics in some parameter space. For the case of a time slice $\mathcal{M}$ in the AdSd+1, all the geodesics can be parameterized by its boundary endpoints $\left(\mathbf{x},\mathbf{y}\right)$, and the Kinematic measure $d\Gamma$ is given by,
$d\Gamma$ $=\text{det}\left(\frac{\partial^{2}\ell\left(\mathbf{x},\mathbf{y}\right)}{\partial\mathbf{x}\partial\mathbf{y}}\right)d\mathbf{x}\wedge d\mathbf{y}=\frac{2^{d-1}}{\left|\mathbf{x}-\mathbf{y}\right|^{2d-2}}d\mathbf{x}\wedge d\mathbf{y},$ (5)
where $\ell\left(\mathbf{x},\mathbf{y}\right)$ is the geodesic length of the geodesic connecting $\mathbf{x}$ and $\mathbf{y}$. Notably, the above Kinematic measure for AdS space is proportional to $\mathcal{I}\left(\mathbf{x},\mathbf{y}\right)$ given in (2), hence we can check that (4) exactly aligns with (3).
One profound implication of the above prescription is that, minimizing the area of a surface $\Sigma_{A}$ homologous to any boundary region $A$ is equivalent to minimizing the number of intersections between $\Sigma_{A}$ (see the red line in Fig.1) and the PEE network. This further implies that we can exactly reproduce the Ryu-Takayanagi (RT) formula for holographic entanglement entropy by developing proper tensor network models where the entanglement entropy of a boundary region is captured by the minimal number of intersections between a homologous surface and the PEE network Lin:2023rxc ; Lin:2024dho ; Wen:2024uwr . Briefly speaking, the RT formula for the boundary region $A$ can be reformulated as follows,
$S_{A}=\min_{\Sigma_{A}}\mathcal{N}(\Sigma_{A})=\min_{\Sigma_{A}}\frac{\operatorname{Area}\left(\Sigma_{A}\right)}{4G}=\frac{\operatorname{Area}\left(\gamma_{A}\right)}{4G},$ (6)
where $\gamma_{A}$ is the famous RT surface.
Figure 1: Extracted from Wen:2024uwr . The PEE network as a tessellation of a time slice of AdS3. Here the blue lines represent the PEE threads, the red line is a surface $\Sigma_{A}$ homologous to the boundary region AA.
3 Geometric setup for bulk tensors
In a model of tensor network, a vertex at site $\mathbf{z}$ with a set of legs corresponds to a quantum state represented by a local tensor $\mathcal{T}_{u_{1}u_{2}\cdots}(\mathbf{z})$, where the index $u_{i}$ characterizes the leg that is associated to a Hilbert space $\mathcal{H}_{i}$ spanned by the basis vectors $\{|u_{i}\rangle\}$. Then the quantum state in the Hilbert space $\bigotimes_{i}\mathcal{H}_{i}$ settled at $\mathbf{z}$ is defined by (assuming Einstein summation hereafter),
$|\mathcal{T}(\mathbf{z})\rangle\equiv\mathcal{T}_{u_{1}u_{2}\cdots}(\mathbf{z})|u_{1}\rangle|u_{2}\rangle\cdots.$ (7)
For simplicity, we set each of $\{\mathcal{H}_{i}\}$ to be the Hilbert space of a $v$-dimensional qudit, hence $u_{i}={1,2,\cdots,v}$.
$\mathbf{z}$$\mathbf{z^{\prime}}$$|\mathbf{z}\mathbf{z^{\prime}}\rangle$
Figure 2: A tensor state located at the bulk site $\mathbf{z}$ (hollow white dot in the center). The legs connecting $\mathbf{z}$ is actually continuously distributed, which can be described by a vector field as in Lin:2023rxc . Only six legs are exhibited, while other legs are collectively represented by the blue region. The contraction between tensors at $\mathbf{z}$ and at adjacent site $\mathbf{z^{\prime}}$ is realized by projecting the total state into a maximally entangled state $|\mathbf{z}\mathbf{z^{\prime}}\rangle$ along the PEE thread in between.
Now we develop the PEE tensor networks. Consider any site $\mathbf{z}$ in a time slice $\mathcal{M}$ of the bulk space, for any PEE thread emanating from it to the boundary, we assign one leg that lies along this PEE thread, hence build a one-to-one correspondence between the PEE threads and the legs connecting to $\mathbf{z}$ (see Fig.2). In the case of Poincaré AdS, any PEE thread emanating from $\mathbf{z}$ intersects the boundary with a unique boundary site $\mathbf{x}$. We denote this thread as $C_{\mathbf{z}}(\mathbf{x})$, and $\mathbf{x}$ can be used to parameterize all the threads or legs associated to $\mathbf{z}$. Let us denote the Hilbert space and basis vectors on each leg as $\mathcal{H}_{\mathbf{x}}$ and $|k_{\mathbf{x}}\rangle$, such that the state (7) is rewritten as
$|\mathcal{T}(\mathbf{z})\rangle\equiv\mathcal{T}_{\mathbf{k}}(\mathbf{z})\left(\bigotimes_{\mathbf{x}\in bdy}|k_{\mathbf{x}}\rangle\right),$ (8)
where $\mathbf{k}$ denotes a collection of indices $\{k_{\mathbf{x}}\}$, and $\{|k_{\mathbf{x}}\rangle\}$ forms an orthonormal basis for the Hilbert space associated to the index $k_{\mathbf{x}}$.
Given any adjacent site $\mathbf{z^{\prime}}$ of $\mathbf{z}$, there is a unique PEE thread connecting them, which picks up a pair of legs associated to $\mathbf{z}$ and $\mathbf{z^{\prime}}$ that lie within this thread (see Fig.2). Then we contract all such pairs of legs to glue adjacent sites along the PEE threads, and finally get a quantum state $|\Psi\rangle$ for the boundary open (uncontracted) legs,
$|\Psi\rangle=\left(\bigotimes_{\mathbf{z},\mathbf{z^{\prime}}\in\mathcal{M}}\langle\mathbf{z}\mathbf{z^{\prime}}|\right)\left(\bigotimes_{\mathbf{z}\in\mathcal{M}}|\mathcal{T}(\mathbf{z})\rangle\right)$ (9)
where $|\mathbf{z}\mathbf{z^{\prime}}\rangle=\frac{1}{\sqrt{v}}\sum_{k=1}^{k=v}|k\rangle_{\mathbf{z}}|k\rangle_{\mathbf{z^{\prime}}}$ is a maximally entangled state defined on any pair of legs connecting the nearby sites $\mathbf{z}$ and $\mathbf{z^{\prime}}$. These contractions connect all the legs into complete PEE threads (see Fig.3) and eventually we get the PEE tensor networks.
$\mathbf{x}$$\mathbf{y}\uparrow\uparrow$$|a\rangle_{\mathbf{x}}$$|b\rangle_{\mathbf{y}}$$\sim T_{ab}(\mathbf{x},\mathbf{y})$$\mathbf{z}_{1}$$\mathbf{z}_{n}$$\mathbf{z}_{2}$$\mathbf{z}_{n-1}\cdots$
Figure 3: Contraction of bulk tensors along a PEE thread results in a PEE thread with two open legs of the boundary sites tangent to the thread. In the factorized PEE tensor network, the thread represents a two-qudit state $T_{ab}(\mathbf{x},\mathbf{y})|a\rangle_{\mathbf{x}}|b\rangle_{\mathbf{y}}$.
4 Factorized PEE tensor network
Let us consider a pair of PEE threads which makes a complete geodesic passing through $\mathbf{z}$, and we denote it as $C_{\mathbf{z}}(\mathbf{x},\mathbf{x^{\prime}})\equiv C_{\mathbf{z}}(\mathbf{x})\cup C_{\mathbf{z}}(\mathbf{x^{\prime}})$. Correspondingly we can divide the legs of $\mathbf{z}$ into pairs and rewrite the quantum state (8) as,
$|\mathcal{T}(\mathbf{z})\rangle=\mathcal{T}_{\mathbf{k}\mathbf{k^{\prime}}}(\mathbf{z})\left(\bigotimes_{\mathcal{C}_{\mathbf{z}}(\mathbf{x},\mathbf{x^{\prime}})}|k_{\mathbf{x}}\rangle|k_{\mathbf{x^{\prime}}}\rangle\right),$ (10)
where $\mathbf{k}$ ($\mathbf{k^{\prime}}$) denotes a collection of indices $\{k_{\mathbf{x}}\}$ ($\{k_{\mathbf{x^{\prime}}}\}$). We define the factorized PEE tensor network in the following:
  • Firstly, we set the tensor $\mathcal{T}_{\mathbf{k}\mathbf{k^{\prime}}}(\mathbf{z})$ to be a tensor product of two-index tensors, each of which is associated to a pair of legs lying within the same complete PEE thread, such that the quantum state (8) becomes a product of states associated to each pair of legs:
    $|\mathcal{T}(\mathbf{z})\rangle=\bigotimes_{\mathcal{C}_{\mathbf{z}}(\mathbf{x},\mathbf{x^{\prime}})}\mathcal{T}_{ab}(\mathbf{z})|k_{\mathbf{x}}\rangle|k_{\mathbf{x^{\prime}}}\rangle\,.$ (11)
  • Furthermore, we require the state $\mathcal{T}_{ab}(\mathbf{z})|k_{\mathbf{x}}\rangle|k_{\mathbf{x^{\prime}}}\rangle$ to be a pair of maximally entangled qudits, i.e. EPR pairs.
Then we contract all the internal legs to get the tensor network. Note that only the legs of adjacent sites are contracted along the PEE thread in between, hence the contractions are always along the PEE threads. More explicitly, let us consider a PEE thread $\mathcal{C}(\mathbf{x},\mathbf{y})$ connecting boundary sites $\mathbf{x}$ and $\mathbf{y}$, and write $\mathcal{C}(\mathbf{x},\mathbf{y})=\{\mathbf{x},\mathbf{z}_{1},\dots,\mathbf{z}_{n},\mathbf{y}\}$ as a set of consecutive adjacent bulk sites with $n\rightarrow\infty$, we contract all the two-index tensors on $\mathcal{C}(\mathbf{x},\mathbf{y})$ to get,
$T_{ab}(\mathbf{x},\mathbf{y})\equiv\mathcal{T}_{ac_{1}}(\mathbf{x})\mathcal{T}_{c_{1}c_{2}}(\mathbf{z}_{1})\dots\mathcal{T}_{c_{n}c_{n+1}}(\mathbf{z}_{n})\mathcal{T}_{c_{n+1}b}(\mathbf{y}),$ (12)
which is also a two-index tensor representing a pair of maximally entangled qudits. The contraction gives us a complete PEE thread with the two open legs tangent to $\mathcal{C}(\mathbf{x},\mathbf{y})$ in the EPR state $|\psi\rangle_{\mathbf{x}\mathbf{y}}=T_{ab}(\mathbf{x},\mathbf{y})|a\rangle_{\mathbf{x}}|b\rangle_{\mathbf{y}}$, (see Fig.3). After the contraction along all the PEE threads, we get a tensor network representing the quantum state for all the boundary open legs:
$|\Psi\rangle=\bigotimes_{{}_{\mathcal{C}(\mathbf{x},\mathbf{y})}}T_{ab}(\mathbf{x},\mathbf{y})|a\rangle_{\mathbf{x}}|b\rangle_{\mathbf{y}}\,$, (13)
which is a product state of EPR pairs. The reduced density matrix for any one of the qudits in the EPR pairs satisfies:
$\rho_{\mathbf{x}}=$ $T_{ab}(\mathbf{x},\mathbf{y})T^{*}_{cb}(\mathbf{x},\mathbf{y})\left(|a\rangle\langle c|\right)_{\mathbf{x}}=\frac{\mathbbm{1}_{\mathbf{x}}}{v},$ (14)
which demonstrates the unitarity of $T_{ab}(\mathbf{x},\mathbf{y})$.
Now we consider any boundary region $A$ in the state (13), and compute the entanglement entropy for $A$. It is convenient to classify all the PEE threads into three classes: (a) $\mathcal{C}^{A}$: threads that anchor only on $A$; (b) $\mathcal{C}^{\bar{A}}$: threads that anchor only on $\bar{A}$; (c) $\mathcal{C}^{A\bar{A}}$: threads that anchor on both $A$ and $\bar{A}$ at each end. Let us rewrite $|\Psi\rangle=|T^{A}\rangle\otimes|T^{\bar{A}}\rangle\otimes|T^{A\bar{A}}\rangle$ where
$|T^{A}\rangle$ $\equiv\bigotimes_{\mathcal{C}(\mathbf{x},\mathbf{y})\in\mathcal{C}^{A}}T^{(A)}_{ab}(\mathbf{x},\mathbf{y})|a\rangle_{\mathbf{x}}|b\rangle_{\mathbf{y}},$ (15)
$|T^{\bar{A}}\rangle$ $\equiv\bigotimes_{\mathcal{C}(\mathbf{x},\mathbf{y})\in\mathcal{C}^{\bar{A}}}T^{(\bar{A})}_{ab}(\mathbf{x},\mathbf{y})|a\rangle_{\mathbf{x}}|b\rangle_{\mathbf{y}},$
$|T^{A\bar{A}}\rangle$ $\equiv\bigotimes_{\mathcal{C}(\mathbf{x},\mathbf{y})\in\mathcal{C}^{A\bar{A}}}T^{(A\bar{A})}_{ab}(\mathbf{x},\mathbf{y})|a\rangle_{\mathbf{x}}|b\rangle_{\mathbf{y}}.$
[…]
$\rho_{A}=\text{Tr}_{\bar{A}}\,\rho=\rho^{(A)}_{A}\otimes\rho^{(\bar{A})}_{A}\otimes\rho^{(A\bar{A})}_{A}\,$. (16)
[…]
$\rho^{(A\bar{A})}_{A}$= $\bigotimes_{\mathcal{C}(\mathbf{x},\mathbf{y})\in\mathcal{C}^{A\bar{A}}}\left(T(\mathbf{x},\mathbf{y})T^{\dagger}(\mathbf{x},\mathbf{y})\right)$ (17)
$=$ $\bigotimes_{\mathbf{x}\in A_{1}}\frac{\mathbbm{1}_{\mathbf{x}}}{v}=\frac{\mathbbm{1}_{A_{1}}}{v^{\mathcal{I}(A,\bar{A})}},$ (18)
where $A_{1}$ denote the qudits in $A$ that are connected to $\bar{A}$ by the threads in $\mathcal{C}^{A\bar{A}}$, and $\mathcal{I}(A,\bar{A})$ counts the number of PEE threads in $\mathcal{C}^{A\bar{A}}$. Therefore, the entanglement entropy between $A$ and $\bar{A}$ is given by:
$S_{A}=S(\rho^{(A\bar{A})}_{A})=\mathcal{I}(A,\bar{A}),$ (19)
where we have set $\log v=1$ such that each thread gives unit contribution to $S_{A}$.
It has been demonstrated that Lin:2024dho , given a static boundary interval $A$ (or any spherical region in higher dimensions) and the corresponding RT surface $\gamma_{A}$, the PEE threads in $\mathcal{C}^{A\bar{A}}$ are the only threads that intersect with $\gamma_{A}$, and each thread only intersects once. This means the number of intersections between $\gamma_{A}$ and the PEE network is exactly $\mathcal{I}(A,\bar{A})$. Since the area of $\gamma_{A}$ can be computed by counting the number of intersections (3), we have
$S_{A}=\mathcal{I}(A,\bar{A})=\frac{\text{Area}(\gamma_{A})}{4G}\,$, (20)
which reproduces the RT formula for any static boundary single intervals and spherical regions. See the appendix A for more details for the bulk geometric picture, and another derivation for the RT formula by constructing isometric mappings from $\gamma_{A}$ to the boundary.
Nevertheless, for disconnected and non-spherical connected boundary regions, the homologous surfaces $\Sigma_{A}$ that only intersect with the threads in $\mathcal{C}^{A\bar{A}}$ no longer exists. Furthermore, the RT surface $\gamma_{A}$ intersects not only with all the threads in $\mathcal{C}^{A\bar{A}}$, but also with a subset of the threads in $\mathcal{C}^{A}$ or $\mathcal{C}^{\bar{A}}$ Lin:2024dho , which means $\mathcal{I}(A,\bar{A})<\frac{Area(\gamma_{A})}{4G}$. For these reasons we fail to reproduce the RT formula for generic boundary regions in this model.
5 Random PEE tensor network
In the second model, we also start from the PEE network, but set the states $|\mathcal{T}(\mathbf{z})\rangle$ (8) on the bulk vertices to be unit random states following Hayden:2016cfa . We will take the random average of such states when computing the entropy quantities. To be specific, we take an arbitrary reference state $|0_{\mathbf{z}}\rangle$ and define $|\mathcal{T}(\mathbf{z})\rangle=U_{\mathbf{z}}|0_{\mathbf{z}}\rangle$ where $U_{\mathbf{z}}$ is a unitary operator, then the random average is equivalent to an integration over $U_{\mathbf{z}}$ according to the Haar probability measure. Again, after contracting all the internal legs of adjacent bulk sites along all the PEE threads using (9), we obtain the random PEE tensor network representing a boundary state for the boundary open legs.
In Hayden:2016cfa , the authors have shown that in a general graph of network with random tensors on the vertices, if we take the large $v$ limit for the bond dimension of legs, the entanglement entropy for an arbitrary boundary region $A$ is dominant by the area term:
$S_{A}=\min\mathcal{N}(\Sigma_{A})\log v,$ (21)
where $\Sigma_{A}$ is a surface in the network homologous to $A$, and $\mathcal{N}(\Sigma_{A})$ counts the number of contracted tensor legs cut by $\Sigma_{A}$. Note that, the above result applies to general network graphs, including our PEE network. In this case, counting the minimal number of contracted legs cut by $\Sigma_{A}$ is equivalent to computing the area of the minimal homologous surface $\gamma_{A}$, which is just the RT surface. Note that, in order to obtain the result (21), we need to set the contribution $\log v$ to the entropy from each thread to be a large number. Under this setup, we need to divide the density of the PEE threads by an overall constant $\log v$ to keep the boundary PEE unchanged. In other words, the density of intersections between any bulk surface and the PEE network becomes $1/(4G\log v)$, such that $\min\mathcal{N}(\Sigma_{A})=\text{Area}(\gamma_{A})/(4G\log v)$. Hence we reproduced the RT formula,
$S_{A}=\min\mathcal{N}(\Sigma_{A})\log v=\frac{\text{Area}(\gamma_{A})}{4G}$ (22)
We emphasize that the result in (21) applies to generic boundary regions, including disconnected and non-spherical connected regions.
6 Discussions
In this work, we assign quantum states to the vertices in the PEE network of AdS space and build two PEE tensor network models, where the entanglement entropy can be explicitly computed. In our models, counting the minimal number of cuts in the tensor network is exactly computing the area of the RT surface in the semi-classical geometry of gravity. This provides the first example of taking natural continuous limit for a discrete tensor network to perfectly match the geometric background.
In the factorized model, the RT formula is only understood for spherical regions. This is not surprising as in this model all the PEE threads are decoupled from each other, which means the model is oversimplified for AdS/CFT. The factorized model reminds us of the bit thread configurations Freedman:2016zud ; Headrick:2017ucz , and we recommend the readers to consult Lin:2023rxc for their similarities and differences  [^2] In Lin:2023rxc it was demonstrated that bit threads configuration for spherical regions $A$ can be reproduced by the superposition of the vector fields describing the PEE threads emanating from $A$..
More interesting models can be developed by including coupling among the PEE threads. In the appendix B, besides regarding the tensors as tensor product of two-index tensors, we also construct a HaPPY-like Pastawski:2015qua PEE tensor network, where we associate a perfect tensor to each site in the PEE network. The RT formula for connected regions in AdS3 is derived using a similar greedy algorithm developed in Pastawski:2015qua for our HaPPY-like PEE tensor network.
As in Hayden:2016cfa ; Pastawski:2015qua , we can also set open legs for the bulk sites to built a holographic code mapping the bulk states to the boundary states, then we can study the bulk quantum correction to the holographic entanglement entropy and identify the analogue of the quantum extremal surface Engelhardt:2014gca . Perhaps the most interesting future direction is to generalize the PEE tensor network to geometric background beyond the Poincaré AdS space, thus construct a new framework to study holographic correspondence beyond AdS/CFT. This is promising because the Crofton formula works for more generic Riemann manifolds, and we can build the network of all the geodesics in the Kinematic Space. The generalization will be non-trivial, as it is quite possible that the geodesics in the bulk will not always anchor on the boundary, which may correspond to spaces containing black holes or singularities.
Acknowledgements.
The authors are supported by the NSFC Grant No. 12447108 and the Shing-Tung Yau Center of Southeast University. We thank Debarshi Basu for helpful discussions and related collaborations.
Appendix A The RT formula in the factorized PEE tensor network
Before making a comparison between the factorized PEE tensor network with the HaPPY code Pastawski:2015qua , we first introduce some preliminary about isometric tensor for the section to be self-contained.
Definition A.1.
An isometry from $\mathcal{H}_{A}$ to $\mathcal{H}_{B}$ is a linear map $T:\mathcal{H}_{A}\rightarrow\mathcal{H}_{B}$ with the property that it preserves the inner product (i.e. $\langle\psi\vphantom{\phi}|\phi\vphantom{\psi}\rangle=\langle T\psi\vphantom{T\phi}|T\phi\vphantom{T\psi}\rangle,~\forall\psi,\phi\in\mathcal{H}_{A}$), such that
  • $\dim(A)\leq\dim(B)$;
  • $T^{\dagger}T=\mathbbm{1}_{A},~TT^{\dagger}=\Pi_{B}$ where $\Pi_{B}$ is a projector on $Im(V)\subset\mathcal{H}_{B}$.
Note that a unitary map is also an isometry, in the case that $\dim(A)=\dim(B)$. Given a complete orthonormal basis $\{|a\rangle\}$ for $\mathcal{H}_{A}$, and $\{|b\rangle\}$ for $\mathcal{H}_{B}$, we represent an isometry $T$ as a two-index tensor:
$T:|a\rangle\mapsto T_{ba}|b\rangle,$ (23)
Then we have
$T_{ba^{\prime}}^{*}T_{ba}=T_{a^{\prime}b}^{\dagger}T_{ba}=\delta_{a^{\prime}a}$ (24)
which we call an isometric tensor.
Proposition A.2.
Given an isometry $T:\left|a_{1}a_{2}\right\rangle\mapsto T_{ba_{1}a_{2}}|b\rangle$ and if $\mathcal{H}_{A}=\mathcal{H}_{A_{1}}\otimes\mathcal{H}_{A_{2}}$, then there exists an isometry (up to a constant) $\widetilde{T}:\mathcal{H}_{A_{2}}\rightarrow\mathcal{H}_{B}\otimes\mathcal{H}_{A_{1}}$ acting as
$\widetilde{T}:\left|a_{2}\right\rangle\mapsto T_{ba_{1}a_{2}}\left|ba_{1}\right\rangle$ (25)
which obeys $\widetilde{T}^{\dagger}\widetilde{T}=\operatorname{dim}\left(A_{1}\right)I_{A_{1}}$.
Proof.
$T^{*}_{ba^{\prime}_{1}a^{\prime}_{2}}T_{ba_{1}a_{2}}=\delta_{a^{\prime}_{1}a^{\prime}_{2},a_{1}a_{2}}=\delta_{a^{\prime}_{1}a_{1}}\delta_{a^{\prime}_{2}a_{2}}$ (26)
$\Rightarrow$ $T_{ba_{1}a^{\prime}_{2}}^{*}T_{ba_{1}a_{2}}=\delta_{a_{1}a_{1}}\delta_{a_{2}^{\prime}a_{2}}=\dim(A_{1})\delta_{a^{\prime}_{1}a_{1}}.$
Proposition A.3.
Given an isometry $T:|a\rangle\mapsto T_{ba}|b\rangle$ and, $\{|a\rangle\}$ and $\{|b\rangle\}$ form complete orthonormal bases for $\mathcal{H}_{A}$ and $\mathcal{H}_{B}$ respectively, then
$\{|T_{a}\rangle\equiv T|a\rangle=T_{ba}|b\rangle:1\leq a\leq\dim(A)\}$ (27)
forms a set of orthonormal vectors for $\mathcal{H}_{B}$. Note that the index $a$ ranges over $\dim(A)(\leq\dim(B))$ values, so that (27) may not forms a basis for $\mathcal{H}_{B}$.
Proof.
[…]
Then let us look at the building blocks of our factorized PEE tensor network, which are the two-index tensors $T_{ab}$ representing pairs of maximally entangled qudits. It implies that each tensor $T_{ab}$ correspond to a unitary map from one leg to the other. Contracting these two-index tensors along PEE threads results in geodesic chords with two-index tensors representing unitary maps from the qudit in one of the endpoints to another. This includes the two-index tensor $T(\mathbf{x},\mathbf{y})$ in (14) for a complete PEE thread. More explicitly, (14) implies
$T(\mathbf{x},\mathbf{y})T^{\dagger}(\mathbf{x},\mathbf{y})=\frac{\mathbbm{1}_{\mathbf{x}}}{v},\quad T^{\dagger}(\mathbf{x},\mathbf{y})T(\mathbf{x},\mathbf{y})=\frac{\mathbbm{1}_{\mathbf{y}}}{v},$ (29)
hence $T(\mathbf{x},\mathbf{y}):|b\rangle_{\mathbf{y}}\mapsto T_{ab}(\mathbf{x},\mathbf{y})|a\rangle_{\mathbf{x}}$ is a unitary map between the pair of qudits shown in Fig.3 at the two boundary sites $\mathbf{x}$ and $\mathbf{y}$. We can regard one leg of $T_{ab}(\mathbf{x},\mathbf{y})$ as input while the other as output, hence define a unitary flow direction for all the threads, i.e. each PEE thread $\mathcal{C}(\mathbf{x},\mathbf{y})$ has one incoming leg and one outgoing leg, giving a unitary map individually.
Now we consider any spherical boundary region $A$ and, as in the main text, classify the PEE threads into three types: the $\mathcal{C}^{A}$, $\mathcal{C}^{\bar{A}}$ and $\mathcal{C}^{A\bar{A}}$, which are represented by the gray, orange and blue semi-circles in Fig.4a respectively for the special case of AdS3. Note that all the PEE threads are semi-circles and the RT surface $\gamma_{A}$ of $A$ is a semi-sphere, it is easy to observe that, all the PEE threads in $\mathcal{C}^{A}$ and $\mathcal{C}^{\bar{A}}$ do not intersect with $\gamma_{A}$, while any thread in $\mathcal{C}^{A\bar{A}}$ intersects $\gamma_{A}$ exactly once. Next we cut the AdS space open along the RT surface $\gamma_{A}$, giving open legs living on $\gamma_{A}$. All the legs on $\gamma_{A}$ are all connected to a subset of the legs on $A$, which we call $A_{1}$. It is convenient to divide the rest of the legs in $A$ into two halves $A_{2}\cup A_{3}$ such that the gray threads connect the legs in $A_{2}$ to those in $A_{3}$. We further set the unitary flow direction for the threads $\mathcal{C}^{A\bar{A}}$ to be from the legs in $\gamma_{A}$ to those in $A_{1}$, and set the flow for the $\mathcal{C}^{A}$ threads to be from $A_{2}$ to $A_{3}$, then the threads in the entanglement wedge of $A$ defines a unitary evolution from the legs in $\gamma_{A}\cup A_{3}$ to the legs $A_{1}\cup A_{2}$. Similarly, we can divide the legs in $\bar{A}$ into $\bar{A}=\bar{A}_{1}\cup\bar{A}_{2}\cup\bar{A}_{3}$, and set $\bar{A}_{1}$ to be the legs connected to $\gamma_{A}$, then the PEE threads defines a unitary evolution from the legs in $\bar{A}_{1}\cup\bar{A}_{2}$ to those in $\gamma_{A}\cup\bar{A}_{3}$. The whole PEE tensor network is equivalent to a quantum circuit shown in Fig.4b.
It is easy to see that, the open legs $A_{2}\cup A_{3}$ do not contribute to the entanglement entropy $S_{A}$, as they just correspond to a set of maximally entangled qudits which make a pure state. The only non-zero contribution to $S_{A}$ comes from the legs $A_{1}$, which are qudits maximally entangled to the legs $\bar{A}_{1}$. If we set the entanglement contained in each thread to be one, then $S_{A}$ is just given by the number of threads in $\mathcal{C}^{A\bar{A}}$, which equals to the number of intersections between $\gamma_{A}$ and the PEE threads. Since the density of intersections on $\gamma_{A}$ is $1/(4G)$ everywhere on $\gamma_{A}$, we reproduce the RT formula $S_{A}=\frac{\text{Area}(\gamma_{A})}{4G}$.
$A\gamma_{A}$
$A_{1}$$A_{2}$$\bar{A}_{3}$$\bar{A}_{2}$$A_{3}$$\bar{A}_{1}$$\gamma_{A}$
Figure 4: (a) The three type of PEE threads, denoted by $\mathcal{C}^{A}$, $\mathcal{C}^{\bar{A}}$ and $\mathcal{C}^{A\bar{A}}$, are labeled by gray, orange and blue solid simi-circles respectively, while the RT surface $\gamma_{A}$ is drawn as the red dashed line; (b) The circuit interpretation of the factorized PEE tensor network. Each PEE thread individually gives a unitary map, and the RT surface $\gamma_{A}$ only cuts the threads in $\mathcal{C}^{A\bar{A}}$ once. The graph can be compared with Fig.19 of Pastawski:2015qua where the “circuit” is built from the net effect of perfect tensors of certain bulk subregions as unitary gates.
It is also useful to derive the RT formula using the arguments in Pastawski:2015qua for more generic tensor network models. To be specific, let us cut the tensor network open along a homologous surface $\mathcal{C}$ whose boundary matches the boundary of $A$, such that the tensor network is divided into two parts $\mathcal{P}$ and $\mathcal{Q}$, which are maps from $\mathcal{C}$ to $A$ and from $\mathcal{C}$ to $\bar{A}$ respectively. The boundary state can be written as
$|\Psi\rangle=\mathcal{P}_{\mathbf{a}\mathbf{c}}\mathcal{Q}_{\mathbf{b}\mathbf{c}}|\mathbf{a}\rangle_{A}|\mathbf{b}\rangle_{\bar{A}}=|\mathcal{P}_{\mathbf{c}}\rangle_{A}|\mathcal{Q}_{\mathbf{c}}\rangle_{\bar{A}},$ (30)
where $\{|\mathbf{a}\rangle\},~\{|\mathbf{b}\rangle\}$ runs over the complete bases for $A$ and $\bar{A}$, and the index $\mathbf{c}$ collects degrees of freedom associated to $\mathcal{C}$. Tracing $\bar{A}$ we get the reduced density matrix,
[…]
whose corresponding entanglement entropy $S_{A}$ is upper bounded by the dimension of the Hilbert space on $\mathcal{C}$, i.e. the number of cuts $\mathcal{N}(\mathcal{C})$ between $\mathcal{C}$ and the network. Then we get the minimal upper bound by choosing the homologous surface with minimal number of cuts, which we also denote as $\gamma_{A}$, hence
$S_{A}\leq\mathcal{N}(\gamma_{A})\log v.$ (32)
Furthermore, according to the proposition (A.3), if the tensors $\mathcal{P}$ and $\mathcal{Q}$ are isometries, then $|\mathcal{Q}_{\mathbf{c}}\rangle$ and $|\mathcal{P}_{\mathbf{c}}\rangle$ are two sets of orthonormal vectors. This implies $\rho_{A}$ is proportional to the identity such that the above inequality is saturated.
In the case of our factorized PEE tensor network model, we have:
  • the homologous surface with minimal number of cuts is exactly the RT surface $\gamma_{A}$, and $\mathcal{N}(\gamma_{A})={\text{Area}(\gamma_{A})}/{4G}$,
  • by regarding $\{\gamma_{A},A_{3},A_{1}\cup A_{2}\}$ as $\{a_{2},a_{1},b\}$ in the proposition A.2, we can construct an isometry $\mathcal{P}$ that maps $\gamma_{A}$ to $A$, likewise we can construct an isometry $\mathcal{Q}$ from $\gamma_{A}$ to $\bar{A}$.
Therefore, the inequality is saturated and the RT formula follows upon setting $\log v=1$.
Appendix B HaPPY-like PEE tensor network
Besides regarding the tensors as tensor product of two-index tensors, we can also construct a HaPPY-like PEE tensor network, where we associate a perfect tensor on each site in the PEE network. Firstly, let us recall the definition of perfect tensor as follows:
Definition B.1.
A $2n$-index tensor $T_{a_{1}a_{2}\dots a_{2n}}$ is a perfect tensor if, for any bipartition of its legs into a set $D$ and complementary set $D^{c}$ with $\left|D\right|\leq\left|D^{c}\right|$, $T$ is proportional to an isometric tensor from $D$ to $D^{c}$, where $\left|\cdot\right|$ denotes the number of legs in this set.
The definition implies that if we assign a flow configuration to a perfect tensor, the tensor always represents an isometric map as long as the number of incoming legs is less than or equal to the number of outgoing legs.
Figure 5: Three local configurations for the legs emanating from one bulk site (the hollow circle) on the homologous surface $\Sigma_{A}$ (red dashed line) in the PEE network. Here, $\mathcal{W}_{\Sigma_{A}}$ is marked as the blue region, and the PEE thread tangent to $\Sigma_{A}$ at this bulk site is marked as the black dashed line.
As was shown in Pastawski:2015qua , a bulk tensor network built from perfect tensors also satisfy the entropy bound (32), and the inequality is saturated if and only if the two tensors $\mathcal{P}$ and $\mathcal{Q}$ obtained by cutting the whole bulk tensor along $\gamma_{A}$ are isometric. Since the graph of our network is the PEE network, again we have $\mathcal{N}(\gamma_{A})={\text{Area}(\gamma_{A})}/{4G}$. In order to reproduce the RT formula, we only need to demonstrate that the mapping $\mathcal{P}$ is an isometry that maps $\gamma_{A}$ to $A$ and $\mathcal{Q}$ is an isometry that map $\gamma_{A}$ to $\bar{A}$.
Now we build the isometric map from $\gamma_{A}$ to $A$ for our HaPPY-like PEE tensor network using an analogous greedy algorithm proposed in Pastawski:2015qua . The procedure is decomposed into the following steps:
  1. 1.
    Given any connected boundary region $A$, let us consider a homologous surface $\Sigma_{A}$ which cuts the PEE tensor network open, and we denote the region enclosed by $A$ and $\Sigma_{A}$ as $\mathcal{W}_{\Sigma_{A}}$ (we do not include the sites on $\Sigma_{A}$ into $\mathcal{W}_{\Sigma_{A}}$). In Fig.5, we mark $\mathcal{W}_{\Sigma_{A}}$ as the blue region and $\Sigma_{A}$ as the dashed red line. Note that, we require the tensor in $\mathcal{W}_{\Sigma_{A}}$ to be an isometric map from the legs on $\Sigma_{A}$ to those on $A$. More explicitly, here the legs on $\Sigma_{A}$ means the legs emanating from $\Sigma_{A}$ into the blue region.
  2. 2.
    Locally we consider any site on $\Sigma_{A}$ and analyze whether the tensor on this site represents an isometry from the legs outside $\mathcal{W}_{\Sigma_{A}}$ to those inside $\mathcal{W}_{\Sigma_{A}}$. In Fig.5, the black dashed line is the PEE thread tangent to $\Sigma_{A}$. Excluding the two legs on the dashed black thread, we find half of the legs stays inside $\mathcal{W}_{\Sigma_{A}}$, while the other half stays outside.
  3. 3.
    If the two dashed legs also stay in $\mathcal{W}_{\Sigma_{A}}$, then the number of legs in $\mathcal{W}_{\Sigma_{A}}$ is greater than the number of outside legs (see Fig.5a), hence the perfect tensor on this site represents an isometric map from the outside legs to the inside legs. In this case we extend $\mathcal{W}_{\Sigma_{A}}$ to swallow this site, hence the extended region still corresponds to a isometric map from the new $\Sigma_{A}$ to $A$.
  4. 4.
    On the other hand, if the two dashed legs stay outside $\mathcal{W}_{\Sigma_{A}}$ (see Fig.5c), then the tensor is not an isometric map from the outside legs to the inside legs, hence we should not absorb the site into $\mathcal{W}_{\Sigma_{A}}$.
  5. 5.
    Therefore, the critical configuration is that $\Sigma_{A}$ locally matches the dashed geodesic (see Fig.5b).
  6. 6.
    Let us start from the $\Sigma_{A}$ as the boundary interval $A$. Obviously it represents an isometry from the open legs to those stretch inside the AdS space. Then $\Sigma_{A}$ moves into the bulk by swallowing more and more bulk sites, until each site of $\Sigma_{A}$ matches the local critical configuration. This means the greedy algorithm halts when $\Sigma_{A}$ matches the RT surface $\gamma_{A}$, such that we arrive at the least upper bound. Furthermore, the entanglement wedge $\mathcal{W}_{\gamma_{A}}$ defines a isometric map from $\gamma_{A}$ to $A$ hence the upper bound (32) is saturated.
The advantage of our HaPPY-like PEE tensor network is that, the greedy algorithm is performed exactly in the AdS space instead of some discrete graph, and the minimal homologous surface exactly matches the RT surface. Nevertheless, as in the HaPPY code, the above arguments only works for connected regions and for a time slice of AdS3.